# Comment on the ”Electric Power Generation from Earth’s Rotation through its Own Magnetic Field”

### v1

@@ Abstract
The suggestion made by C. F. Chyba and K. P. Hand about electric power generation from Earth’s rotation through its own magnetic field is intriguing [1, 2]. Here, we reanalyze the theoretical arguments for their conclusion, especially those presented in [1]. The model that they consider, is a magnetic cylindrical shell moving with low velocity $v$ in the $y$ direction at a right angle to the direction of the Earth’s magnetic field $B_{\infty}$. We first focus on one of the electromagnetic boundary conditions for moving dielectric boundaries (a topic not dealt with in [1, 2]), and obtain therefrom a result at variance with the main result in the mentioned papers. Therefore, we have to conclude that their result does not satisfy this boundary condition. Moreover, by a simple calculation, we observe that a displacement of the shell in the mentioned $y$ direction, does not involve any expenditure of external work. A process involving dissipation of energy as envisaged in [1, 2] thus appears to be forbidden for energy conservation reasons.
@@ I Introduction
There have appeared some criticisms of this idea - cf. for instance the paper of Jeener [4] and its rebuttal in [5]. It is also notable that in an earlier experiment of Veltkamp and Wijngaarden no energy production of this magnitude was found [6]. As expressed by Wijngaarden in [3]: ”I am still convinced that the theory of Chyba et al. cannot be correct”.
In view of the uncertainty of the situation, we decided to make a reconsideration of the theory presented in [1] and [2], although their theory is detailed and very carefully done. Mainly, we will focus on their first paper [1]. Actually, we find that there is a missing piece in the formalism, namely the electromagnetic boundary condition on dielectric surfaces in motion. We discuss the implications of this omission for the main conclusion of the whole analysis, as our results are at variance with each other. We also find that the principle of energy conservation gives an important restriction in this case.
@@ II The case $\bf v=0$
The geometry of the setup is clearly sketched in [1] and [2], and will not be reproduced here. It will however be convenient to reproduce the expressions for the components of the static magnetic field, called ${\bf B}_{0}$, both outside and inside the dielectric shell, in terms of cylindrical coordinates $\{\rho,\phi,z\}$,
@@
Here $y=\rho\sin\phi$, and the expressions for the coefficients are
@@ III The case $\bf v\neq 0$}
When the shell moves with constant velocity $\bf v$ in the $y$ direction, the Maxwell equations within the shell will be modified. We call the moving frame of reference $K$, while the static reference system considered in the previous system is called $K_{0}$. Since the moving case involves energy dissipation, the conductivity $\sigma$ of the material will have to be accounted for. In analogy to the well known Reynolds number in hydrodynamics, in Refs. [1, 2] was defined a magnetic Reynolds number, $R_{m}=v\xi/(\sigma\mu)^{-1}$, where $\xi$ is a typical length over which the current flow changes appreciably, and $(\sigma\mu)^{-1}$ is the analog of the hydrodynamic kinematic viscosity. It is natural to associate $\xi$ with the outer radius $b$.
@@
and $A_{t}(\rho,\phi,t)$ relates to the time dependent equation
@@
where $K_{1}$ is a Bessel function. From this the gradient of the scalar potential is derived as $\nabla V=-v\beta_{1}\hat{\bf z}$. Correspondingly, the axial electric field in the interior of the shell, i.e. ${\bf E}=-\nabla V-\partial{\bf A}/\partial t$ becomes
@@
when omitting the last term. The electric field on the outside of the shell is obtained by setting $\mu_{r}=1$, whereby $\beta_{1}=B_{\infty}$ and so
@@
There are two boundary conditions for dielectric boundaries moving with small velocities [7]. When applied to $\rho=b$ they read
@@
where $v_{n}$ is the velocity of the boundary along the normal $\bf n$. We shall focus only on Eq. (22). Inserting the expression above for $\bf E$ on the left hand side, we first get
@@
We want to solve this equation with respect to the component $B_{x}(b-)$ on the inside surface. For that we need the expression for the exterior component $B_{x}(b+)$ also. This component was not given explicitly in [1, 2], but can be derived as follows, making use of the same method. We start from Eq. (11), modifying the last term, $A_{0}(\rho>b)=B_{\infty}y+h(z)$, where $h(z)$ is determined from Eq. (16) to be $h^{\prime\prime}(z)=2kB_{\infty},$ or $h(z)=kB_{\infty}z^{2}.$ Going to the case $v\neq 0$, we get the extended form
@@
with $C$ a constant, and $K_{1}(k\rho)=1/(k\rho)$ approximately for small values of $k\rho$. The value of $C$ is determined from requiring agreement with the expression (11) for $v=0$ for small values of $ky$, implying $C=\beta_{3}kb^{2}$, leading to the external potential
@@
is the magnetic Reynolds number. In a vacuum, $\sigma=0$, so we can just identify $A_{z}$ with $A_{0}$ on the outside.
Actually, this could have been seen in a more direct way, by exploiting the gauge condition $\nabla\cdot A=0$, which was found to apply also when $v\neq 0$ (cf. Eq. (22) in [1] in the vacuum case $\eta\rightarrow\infty$.
@@
Thus the final result, in the same frame $K$, becomes
@@
Thus, our application of the electrodynamic boundary conditions for moving boundaries leading to the expression (31), does not agree with the expression (32) from Ref. [1]. Although this fact is disturbing, it does not in opinion solve the essence of the problem. The main point is rather that a process as envisaged herein cannot occur at all, due to energy conservation. This will be elaborated upon in the next section.
@@
Therefore, our inclusion of the electromagnetic boundary conditions for moving boundaries leading to the expression (31), disagrees with the expression (32) given in Ref. [1].
IV Discussion and Summary
We go back to the static case $v=0$, and calculate the total magnetic force component $F_{y}$ on the shell, per unit length in the $z$ direction. To this end we only need the Maxwell stress tensor $T_{ik}$ on the outside, at $\rho=b+$. An important point is that the surface stress originates from the volume force density ${\bf f}=-(1/2)H^{2}\nabla\mu$ acting in the interior of the dielectric boundary layer. Upon integrating the normal component of this force across the boundary region one obtains the surface pressure in its conventional form, as a difference between the normal components of the Maxwell tensor on the two sides of the surface. This force always acts towards the optically thinner region, which in our case with $\mu_{r}>1$ means that it acts outwards (cf., for instance, [8, 9, 10]).
For $\rho>b$, the only component of the Maxwell tensor needed is
$T_{\rho\rho}=\frac{1}{\mu_{0}}B_{0\rho}^{2}-\frac{1}{2\mu_{0}}{\bf B}_{0}^{2}.$ (34)
From Eqs. (1) and (2) we have on the outer surface
$B_{0\rho}(\rho+)=(B_{\infty}+\beta_{3})\cos\phi,$ (35)
${\bf B}_{0}^{2}(\rho+)=B_{\infty}^{2}+2B_{\infty}\beta_{3}\cos 2\phi+\beta_{3}^{2}.$ (36)
Inserting these expressions into the integral we obtain
$F_{y}=\int_{0}^{2\pi}T_{\rho\rho}(b+)\rho\sin\phi d\phi=0.$ (37)
We may now present the following physical argument: In the frame $K_{0}$, let the shell be displaced by an amount $\Delta y$ in the $y$ direction. The mechanical work spent is thus $\Delta W=F_{y}\cdot\Delta y=0$. But it is impossible that such a process where no work has been done, can lead to an electric current, producing energy. According to this, we must conclude that the process envisaged in [1, 2] cannot work. The same can be said for a stationary situation, in which the shell moves with constant velocity $v$ for an arbitrary time.
Based on our analysis, one can conclude that the process of electric power generation from Earth’s rotation cannot start at all under the given circumstances, contrary to the conclusion made in [1, 2].
It seems to us, however, that if one places the dielectric shell in another direction, which has an angle with respect to the Earth’s magnetic field, one might arrive at a different conclusion, namely to produce electric current. This case is currently under study.
@@ Acknowledgements
We are grateful to Anca Tureanu for illuminating discussions and reading the manuscript.

### v2

@@ Abstract
The suggestion made by C. F. Chyba and K. P. Hand about electric power generation from Earth’s rotation through its own magnetic field is intriguing [1, 2]. Due to the importance of the subject, we have re-analyzed the theoretical arguments and derivations leading to their conclusion, by paying special attention to several issues possibly neglected before. The model they consider is a magnetic cylindrical shell moving with velocity $\mathbf{v}$ in the $y$ direction at a right angle to the direction of the Earth’s magnetic field $\mathbf{B}_{\infty}$. First we analyze the electromagnetic boundary conditions when the shell is moving with a constant velocity $\mathbf{v}$, as this point, although of importance, has not been taken care of in [1, 2]. Indeed, this procedure leads us to differences in the values of electromagnetic fields when compared with the expressions given in the cited references. Second and as a result, we find that the mechanical force created by the moving shell becomes different from the one derived in [1, 2]. Obviously, the expression for the amount of electric power generation from Earth’s rotation will also be different from the previously obtained one. The latter is important for evaluating the amount of produced power, maximizing it by choosing the parameters of the shell, and for the comparison with experimental findings.
@@ I Introduction
The result obtained by Chyba and Hand some years ago about generation of electric power from Earth’s rotation through its own magnetic field has generated a lot of interest [1], even more so after the recent announcement about preliminary experimental results seemingly supporting the theoretical predictions stated in [2]. As surveyed in [3], physicists are divided with respect to the claim.
The geometrical model of the theory as well as the model used in the experiments is that of a conducting cylindrical long shell of inner radius $a$ and outer radius $b$, where the material in the region $a\leq\rho\leq b$ is magnetic, with permeability $\mu$. In the theoretical description of the system, a frame of reference is taken, where the shell is moving with a constant velocity $\mathbf{v}=v\hat{\mathbf{y}}$, $\mathbf{v}$ being the Earth rotation velocity at the place of the shell and considered to be linear, in a direction transverse to the $x$ direction of the static magnetic field of the Earth, called $\mathbf{B}_{\infty}$. Thus, $\mathbf{B}_{\infty}=B_{\infty}\hat{\mathbf{x}}$. $\hat{\mathbf{x}}$ and $\hat{\mathbf{y}}$ are the unit vectors along the axes $x$ and $y$. The slight declination, at present 9.2 to 11.3 degrees, of the magnetic field relative to the geographic axis has been justifiably neglected in the estimates.
Some criticisms of this idea have appeared, cf., for instance, the paper of Jeener [4] and its rebuttal in [5]. It is also notable that in an earlier experiment of Veltkamp and Wijngaarden, no energy production of this magnitude was found [6]. As expressed by Wijngaarden in [3]: ”I am still convinced that the theory of Chyba et al. cannot be correct”.
In view of the apparent uncertainty of the situation we have decided to make a reconsideration of the theory presented in [1] and [2], although their theory is detailed and very carefully done. We shall focus mainly on their first paper [1]. Evidently, there is a missing piece in their treatment, namely the electromagnetic boundary conditions on the shell material surfaces in motion. We discuss the implications of this omission for the main conclusion of the whole analysis, as our results are at variance with the once presented in the cited papers. We find that our expression for mechanical force associated with the motion is different compared with the one obtained in [1] and [2]. Obviously, the mechanical work and produced power will also be different.
@@ II The case $\bf v=0$
The geometry of the setup is clearly sketched in [1] and [2], and will not be reproduced here. It will however be convenient to reproduce the expressions for the components of the static magnetic field, called ${\bf B}_{0}$, both outside and inside the conductive shell, in terms of cylindrical coordinates $\{\rho,\phi,z\}$,
@@
Here, $y=\rho\sin\phi$ and the expressions for the coefficients are
@@ III The case $\bf v\neq 0$}
When the shell moves with constant velocity $\bf v$ in the $y$ direction with respect to our rest frame, the Maxwell equations within the shell will be modified. We shall call this shell-moving frame of reference $K$, while the shell-at-rest reference frame considered in the previous system is called $K_{0}$. Since the moving case involves energy dissipation, the conductivity $\sigma$ of the material will have to be accounted for. In analogy to the well known Reynolds number in hydrodynamics, in Refs. [1, 2] was defined a magnetic Reynolds number, $R_{m}=v\xi/(\sigma\mu)^{-1}$, where $\xi$ is a typical length over which the current flow changes appreciably, and $(\sigma\mu)^{-1}$ is the analog of the hydrodynamic kinematic viscosity. It is natural to associate $\xi$ with the outer radius $b$.
@@
and $A_{t}(\rho,\phi,t)$ relates to the time-dependent equation
@@
where $K_{1}$ is a Bessel function. From this the gradient of the scalar potential is derived as $\nabla V=-v\beta_{1}\hat{\bf z}$. Correspondingly, the axial electric field in the interior of the shell, i.e. ${\bf E}=-\nabla V-\partial{\bf A}/\partial t$, becomes
@@
when omitting the last term in (19). The electric field on the outside of the shell is obtained by setting $\mu_{r}=1$, whereby $\beta_{1}=B_{\infty}$, and so
@@
There are two boundary conditions for dielectric boundaries moving with small velocities [7]. When applied to $\rho=b$, they read
@@
where $v_{n}$ is the velocity of the boundary along the normal to the boundary surface $\bf n$.
For a medium with zero conductivity, this boundary condition is straightforward. That the same condition holds also in the case of a general conductivity, when, as usual, the electric charges are distributed over the volume of the material without singular points or regions, follows from the Maxwell equations. First, Eq. (22), related to the rest-frame equation ${\bf\nabla\times E}=-\partial{\bf B}/\partial t$, does not contain $\bf J$ at all. Second, in Eq. (23), related to ${\bf\nabla\times H}={\bf J}+\partial{\bf D}/\partial t$, we can perform an integration over a coin-shaped volume whose main direction aligns with the surface. The volume contains the material surface. Taking the limit when the thickness of the coin goes to zero, we see that the influence from $\bf J$ has to vanish. Only in exceptional cases containing singular currents, for instance, would $\bf J$ be able to contribute.
Inserting the expression (20) and (21) for $\bf E$ into the left-hand side of (22), we get
@@
We should now solve this equation with respect to the component $B_{x}(b-)$ inside the surface. For that we also need the expression for the exterior component $B_{x}(b+)$. This component was not given explicitly in [1, 2], but can be derived as follows, making use of the same method. We start from Eq. (11), modifying the last term, $A_{0}(\rho>b)=B_{\infty}y+h(z)$, where $h(z)$ is determined from Eq. (16) to be $h^{\prime\prime}(z)=2kB_{\infty},$ or $h(z)=kB_{\infty}z^{2}.$ Going to the case $v\neq 0$, we get the extended form
@@
with $C$ a constant, and $K_{1}(k\rho)=1/(k\rho)$ approximately for small values of $k\rho$. The value of $C$ is determined by requiring agreement with the expression (11) for $v=0$ for small values of $ky$, implying $C=\beta_{3}kb^{2}$, leading to the external potential
@@
is the magnetic Reynolds number. In vacuum, $\sigma=0$, so we can just identify $A_{z}$ with $A_{0}$ on the outside.
Actually, this could have been seen in a more direct way, by exploiting the gauge condition $\nabla\cdot A=0$, which was found to apply also when $v\neq 0$ (cf. Eq. (22) in [1] in the vacuum case $\eta\rightarrow\infty$).
@@
Thus, the final result, in the same frame $K$, becomes
@@
Thus, our inclusion of the electromagnetic boundary conditions for moving boundaries leading to the expression (31), disagrees with the expression (32) given in Ref. [1].
IV Produced power
In this section we evaluate the produced power and also the dissipation, although they have previously been dealt with in [1, 2]. Due to our different values for the induced fields as compared with the ones obtained in [1, 2], the expression for the power and dissipation will also be different. The differences between the results obtained in [1, 2] and here all stem from the fact that we have taken into account the requirement of boundary conditions. In the next section we elaborate on the necessity and importance of boundary conditions and illuminate it by mentioning a few known examples.
We consider the system of reference in which the shell is moving.
Assuming that the medium has conductivity, we start from Eq. (20) which shows that in the frame K (in which the shell is moving) there is a constant electric field $\bf E=v\beta_{1}\hat{\bf z}$ in the $z$ direction. There is thus an electrical force density component $\rho\bf E$} in that direction, $\rho$ being the charge density. This force yields no work in the $y$ direction.
Secondly, we consider the magnetic part $\bf J\times B$ of the Lorentz force density, where ${\bf J}=\sigma{\bf E}$ with $\sigma$ the conductivity. What expression is to be inserted for $\bf B$ here? The expression should be chosen such that the force reduces to zero in the static case. So, when subtracting the latter fields, we obtain simply zero for the velocity-induced Lorentz force.
Of main interest is the induced magnetic field ${\bf B}_{\rm induced}$. This field is calculated from a Galilean transformation between the two inertial systems as ${\bf B}_{\rm induced}=(1/c)^{2}{\bf v\times E}$ and lies in the $x$ direction. Thus, ${\bf J\times B}_{\rm induced}=\sigma{\bf E}\times[\frac{1}{c^{2}}{\bf v\times E}]=\sigma\beta E^{2}\hat{\bf y}/c$, where $\beta=v/c$. This force in principle does involve work in the $y$ direction. The corresponding power $P$ per unit volume and time is $P=\sigma\beta E^{2}{\bf v\cdot\hat{y}}/c=\sigma\beta^{2}E^{2}.$ Now, we make use of the order-of-magnitude relation $E\sim Bc$, which follows from Maxwell’s equation ${\bf\nabla\times E}=-\partial{\bf B}/\partial t$ if we insert $|{\bf\nabla\times E}|\sim E/\lambda$, $|\partial{\bf B}/\partial t|\sim\omega B=(2\pi/\lambda)Bc$, with $\lambda$ the wavelength. One can alternatively express the power in terms of the magnetic field as $P=\sigma\beta^{2}B^{2}c^{2}$ (here $B$ meaning the induced field). As $\beta\ll 1$, the factor $\beta^{2}$ makes $P$ far too small. By subtracting the static fields, we obtain an expression close to zero. This is a main result in our investigation.
V Boundary conditions
In our paper, an important point was the application of one of the electromagnetic boundary conditions on dielectric surfaces in motion, using general formulas from Ref. [7]. These are general conditions, that make no restriction on the magnitude of the conductivity $\sigma$. The physics of the boundary conditions is the following: In the inertial frame where the instantaneous velocity of the boundary is zero, the conventional conditions, namely the tangential component of the electric field $\bf E$ as well as the tangential component of the magnetic field $\bf H$, must be continuous. If they were not, this would lead to infinite surface charges or infinite surface currents at the boundary, and that is obviously not the case here. These conditions, as we have mentioned above, are valid irrespective of whether the conductivity $\sigma$ has an extremely low value (as is the case for a dielectric) or is finite for the case of a conductor shell, as far as $\sigma$ is a constant independent of frequency.
To clarify still more the essence of the present Comment, its importance and differences with the original treatment in [1] and [2], let us emphasize the following points:
In classical and also in quantum physics described by differential equations, without imposing boundary conditions one could add several arbitrary (constants and constants multiplied by coordinates to some powers) terms to any solution of the differential equations and in that case obtain any value, including infinity, for the physical quantities of the system under study, e.g. its total energy could become infinite.
A typical example in quantum mechanics is when we consider the solution of Schrödinger’s equation at the boundary of a potential, e.g. a harmonic oscillator; one has to make sure that the wave function itself, as well as its first order derivative, are continuous at the boundary of the potential, which are the matching conditions, otherwise the equation would lose its meaning due to singularity and infinite value of the derivatives.
The solution of Maxwell’s equations, which is the main theoretical basis of the present work to start with, is the same and one should impose the boundary conditions, otherwise the solutions are not unique and the physical quantities can be arbitrary and different. This is the difference between the treatment of the present work and the original one performed in [1] and [2]
In the present work it is shown that the expressions for work performed and for dissipation (although negligible) are different from the expressions obtained in the original treatment performed in [1] and [2]. Since the differences are in the form of different parameters of the shell, they are important for experimental observations and for planning to maximize the amount of electric power generation.
Another difference between the two treatments is that, while we consider the Maxwell equations and their solutions in a certain rest frame, in [1] and [2] they are considered in a moving system. However, it is well known that the physical quantities are invariant and independent on the system of coordinates. Here, the work ${\bf F}\cdot{\bf r}$ is an invariant scalar, since the system in such a nonrelativistic motion of the Earth has symmetry under the Galilean transformations.
Therefore, the source of the discrepancy between the results of the present work and of the ones in [1] and [2], comes only from the imposition of the boundary conditions.
VI Discussion and Summary
For comparison, as already mentioned, the effective value of the product $\bf E\cdot J$} obtained by Chyba and Hand is different from that we have found above. Writing an extra subscript CH to refer to their modified current density, we reproduce from Eq. (81) in [1]:
${\bf E\cdot J}_{CH}=\sigma v^{2}\beta_{1}(\beta_{1}-B_{x}).$ (34)
Here, the static contribution has been subtracted off. The power provided to the shell is found by integrating this expression over the volume.
The last equation (34) for specific energy dissipation enables us to make an important observation. The magnetic field component $B_{x}$, referring to the inertial frame where the shell is moving, is given in the interior of the shell by Eqs. (59) and (60) in Ref. [1], as the difference between the static field component $B_{0x}$ and a component $B_{1x}$. The latter is small, as is seen from the ratio $B_{1x}/B_{0x}\sim R_{m}=\mu\sigma vb$ which contains $v$. Thus, the contribution from $B_{1x}$ to the power is negligible. What remains is the difference $\beta_{1}-B_{0x}$, which is given by Eq. (10a) in the same reference. This difference contains $\cos 2\phi$ : the expressions for $\beta_{1}$ and $B_{0x}$ are given by eqs. (7) and (3), respectively. Upon integration over $\phi$ from 0 to $2\pi$ it vanishes, so that the total power given to the shell will be zero. That is barely an acceptable global result. However, the local dissipation values, containing a cosine function, oscillate and become negative in a large region of $\phi$ around the shell, what is physically not acceptable.
@@ Acknowledgements
We thank Markku Oksanen and Anca Tureanu for illuminating discussions and reading the manuscript, and are grateful to Christopher F. Chyba, Kevin. P. Hand, and Thomas H. Chyba for valuable comments on the first version of this paper.